OpenStudy (anonymous):

Solve tan^2(x) - 3tan(x) + 2=0

OpenStudy (anonymous):

Quadratic in tanx

OpenStudy (anonymous):

let tan(x)=t then the equation becomes: t^2-3t+2=0 (t-2)*(t-1)=0 t=2 or t=1 tan(x)=2 or tan(x)=1(backsubstitution) x=63.43 degrees or x=45 degrees